Factorise. (l+m)^2 -(l-m)^2 ?????????????????????????/

Factorise.
 (l+m)^2 -(l-m)^2 ?????????????????????????/

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(l+m)^2-(l-m)^2
= {(l+m)-(l–m)}{(l +m)+(l–m)}
Using Identity: x^2-y^2= (x+y)(x-y)
= (l+m–l+m)(l+m+l–m)
= (2m)(2l)
= 4 ml

Thanks